OpenStudy (dallaslifebaby01):

Demonstrate how you can use the slope formula to write the point-slope form of an equation of a line.

OpenStudy (candy13106):

do u know the slope formula and point slope form

OpenStudy (dallaslifebaby01):

no i don't

OpenStudy (candy13106):

ok the slope formula is \[y2-y1/x2-x1\]

OpenStudy (candy13106):

thats how u find m then u use m and plug it into the point slope

OpenStudy (candy13106):

and point slope form is \[y-y1=m(x-x1)\]

OpenStudy (dallaslifebaby01):

hey so sorry i didn't respond, my computer crashed @Candy13106

OpenStudy (candy13106):

its ok so u get the y1, x1, and m

OpenStudy (dallaslifebaby01):

yeah i think

OpenStudy (candy13106):

ok

OpenStudy (sooobored):

slope formula \[m=\frac{y_2 - y_1}{x_2-x_1}\] point- slope formula \[(y_2 -y_1) = m (x_2-x_1)\] to go from slope formula, to point-slope multiply both sides of the slope formula by (x2-x1) in order to remove it from the denominator and get a from similar to point-slope

OpenStudy (skullpatrol):

An equation of a line can be found from (a) the slope and the y-intercept, (b) the slope and any point on the line or (c) any two points on the line.

OpenStudy (skullpatrol):

do you want some examples?

OpenStudy (dallaslifebaby01):

that would be helpful

OpenStudy (skullpatrol):

let's start with (a)

OpenStudy (dallaslifebaby01):

okay, the slope is what it changes by right?

OpenStudy (dallaslifebaby01):

like something decreases by 3 the slope would be 3?

OpenStudy (skullpatrol):

yes

OpenStudy (dallaslifebaby01):

okay

OpenStudy (skullpatrol):

but think of it as 3/1 = rise/run

OpenStudy (skullpatrol):

also for "decreases" it would be -3/1

OpenStudy (skullpatrol):

does that^ make sense?

OpenStudy (dallaslifebaby01):

yes that does thank you

OpenStudy (skullpatrol):

have you seen this formula: $$\Huge y=mx+b$$

OpenStudy (dallaslifebaby01):

yes

OpenStudy (skullpatrol):

cool any questions?

OpenStudy (dallaslifebaby01):

no i think i understand all this much bette enow, thank you!

OpenStudy (skullpatrol):

Thanks for trying to understand :-)